OpenStudy (anonymous):

Bob's Car Rental is offering a special of $45 a day for a convertible as long as you purchase the car damage protection insurance for $30. You will need the car for 5 days. How much of your budget is going to rent the car?

OpenStudy (jollypikachu300):

hey @elsaismypuppydog what u do t multiply 45 times 30

OpenStudy (anonymous):

The price is $45 (m) per day and we need the car for 5 (x)days. The insurance is an additional $30 (b). We are trying to find the total cost for the rental (y). y = 45x + 30 when x = 5 y = 45 (5) + 30 y = $255 It costs $255 to rent the convertible for 5 days.

OpenStudy (jollypikachu300):

ok u shoukld have added the anwsers

OpenStudy (jollypikachu300):

it is y = $255 It costs $255 to rent the convertible for 5 days. thats the anwser